# Spectropolarimeter on a 2–4 m class telescope and proposed science cases

### v1

@@ 1 Introduction
Interstellar magnetic fields can be mapped with the polarization of starlight caused by aligned dust grains in different lines-of-sights and in different environments. This was first observed in optical wavelengths and was explained to be arising from dichroic extinction by elongated dust grains aligned with the ambient magnetic field. Interstellar polarization also provides an opportunity to study the dust grain characteristics e.g. size distribution and composition.
An important science to be addressed with a spectropolarimeter is the investigation of interstellar dust. The differential extinction giving rise to the polarization curve (parameterized by the empirical, convex, “Serkowski function”
where is the maximum polarization, located at , and “K” controls the width of the curve. As shown by (ref. 17), the Serkowski form can be understood in terms of Mie scattering off an underlying MRN grain size distribution - if grains smaller than 0.045m are not aligned. The Serkowski fit to the polarization spectrum of star-light shinning through this interstellar dust can therefore constrain the size of aligned dust grains. Adding extinction information to the amount of polarization provides important complementary information that can break the degeneracy between the underlying grain size distribution, and the size dependent grain alignment mechanisms, and their efficiencies. Combining spectroscopy and polarimetery is a time efficient and better approach to perform such investigations in comparison to the often used technique of imaging polarimetry with use of filters.
@@ 2.1 Interstellar Medium Science
Spectropolarimetry of stars in or behind nearby star-forming regions: This type of study is needed to obtain the polarization curve for a number of environments and source types. This provides a direct means to probe the size distribution of ISM dust. It is a powerful technique that for general ISM studies need 2-4 m telescopes, which can use background stars of any spectral class and provides inherently complementary probes to infra-red data. For spatially restricted regions, where the background star selection is severely restricted, requiring a fainter magnitude limit, 4 m telescopes become necessary. As discussed above, the ISM polarization spectrum shows a convex shape with a peak in the range = 0.40.7m. Modern “Radiative Alignment Torque (RAT) alignment” theory, explains this relation by predicting that grains are aligned if exposed to an anisotropic radiation field with wavelengths less than the grain diameter. A universal linear relationship between and extinction () has been found by (ref. 3) .
(2) where is the average total-to-selective extinction. This result can be understood as due to the progressive reddening of the radiation field into the clouds, but with a fixed, underlying, total grain size distribution throughout each cloud. The dependence on indicates that the average grain size varies from cloud to cloud. As discussed below in item 4, grain growth in the denser parts of the cloud can be traced from the deviations of the linear relationship in equation 2.
@@
Investigating grain alignment efficiencies in the regions filled with FUV radiation (high intensity radiation fields) This project is an extension of our study of the grain alignment variations around the wall of the Local Bubble and our recently published paper on IC59 and IC63 clouds illuminated by B0 IVe star Cas. We used our polarization and extinction measurements to estimate polarization efficiencies in these clouds. We investigated how these polarization efficiencies vary with the flux from gamma Cas. Is this UV illumination related to the polarization and hence grain alignment efficiencies in the clouds? This experiment was performed on IC 59 and IC 63 and we compared our results with other regions. We saw a change in the slope of the in polarization efficiencies as we went toward the higher illumination sides. Since IC 59/63 clouds are relatively closer to their illuminating source as opposed to regions like local bubble, we saw high poloarization efficiencies in these clouds. Figure 1 below shows these results. However, as we move towards the regions with more luminous source(s), such as LDN 204 and especially the Orion Veil, we see that the polarization efficiency seem to saturate (Soam & Andersson., in prep.). We speculate that this is either because in the clouds closer to the radiation source all the alignable grains have aligned by the radiation, or that as the radiation field becomes very strong the large grains are destroyed through the so called RATD destruction. Because both RAT alignment and RAD Destruction are more efficient for larger grains, but change the polarization efficiency in opposite directions, broad-band spectro-polarimetry would allow the two possibilities to be separated. Our proposed spectropolarimeter is the best suit for that purpose.
@@
The circumstellar envelopes of asymptotic giant branch (AGB) stars provide unique mineralogical laboratories for studying the effects on the grain alignment. This is because of the dredge-up of newly synthesized carbon during the thermal pulses of the AGB phase and the stability of the CO molecule. When the star first ascends the AGB the circumstellar envelope (CSE) is oxygen rich with all the carbon tied up in CO molecules. The dust is therefore made up of silicate and metal oxide grains. As the dredge-up of newly processed material is transferred into the CSE, the C/O ratio eventually exceeds unity and all the oxygen is tied up in CO, with the dust now made up from carbon materials. Thus, the two types of AGB envelopes provide the opportunity to study the silicate and carbon dust separate from each other. As with interstellar medium studies, the Serkowski-curve parameters can yield important information about the dust in the CSE. Grain formation and its influences on the formation and characteristics of the AGB winds can be studied by probing the light from field stars background to the AGB CSE. Models of the initiation and mass loss characteristics of these winds are still an unresolved issue. The brightest AGB CSEs can be probed by FIR and sub-mm wave data at high angular resolution. However, for such emission polarimetry the complex interaction of dust temperature, emissivity and column density/opacity can be difficult to disentangle. With telescopes and polarimeters capable of reaching V 16 in (at least) BVRI polarization curves at several different points in the CSEs can be acquired, probing the alignment efficiency as well as the grain size distribution (at different radii). In initial studies of the IRC+10216 CSE we see a large and K (although the Serkowski form may not be ideal) indicating large grains as compared to the ISM. This is consistent with the results of ab initio AGB wind models, but better S/N is needed. We note that, in the intense radiation field from the central star, the RATD grain destruction mechanism should be important in these CSEs, Measuring spectral types (and thereby the extinction), and polarization of a number of stars shining through the envelopes of the AGB stars can provide observational constrains on the dust grain size in the AGB stars. The fact that our observations indicate large grains without a small grain population indicates that such destruction is not important. This may indicate grains with high tensile strength and hence compact solid structures. Again more and better data are needed to fully explore these possibilities. Based on FIR/sub-mm wave maps of the nearest AGB stars and Gaia distance and photometry many of the nearby AGB stars have several (in some cases 10) background field stars within the projected 70m extent of their CSEs. An extension of such studies would be to probe the ionized regions of planetary nebulae, where the evolved star and dust interacts with the hard UV of the central star.
@@
Grain Growth and Destruction The polarization curve, under RAT alignment, is determined by the size distribution of aligned grains and the SED of the radiation field. The smallest aligned grain is given by the largest of the smallest grain that satisfies the RAT criterion and smallest grain not disaligned by collisions.
d, where is the shortest wavelength available and d is the grain diameter: Lyman limit (912Å) for the diffuse, neutral, ISM; shorter for fully ionized gas close to hot stars, or longer for material inside of ISM clouds where reddening is active.
Thermal collisions disalign the smallest grains first. The collision rate goes at and the grains disalignment time goes as 1/d. Hence for denser, warmer gas the limit of aligned grains grows. The over-all grain size distribution, and the upper cut-off of it is determined by grain growth and destruction.
In dense gas (clumps in clouds) the density can become large enough that significant grain growth can occur, extending the size distribution and affecting the red part of the polarization curve.
If significant grain destruction (in ISM shocks, or through the RATD mechanism) take place, the largest grains may be destroyed shifting the polarization curve to the blue.
These various effects can be probed by high-quality determination of the polarization curve in different environments. (ref. 41) studied the polarization curve variations in the Taurus molecular cloud and found evidence for the effects of reddening, gas-dust collisions disalignment and grain growth in clumps dark and cold enough for ISM ices to have formed. This is manifested as a bifurcation in the vs. A relation (equ. 2), such that, beyond A mag., a second, much steeper relationship was added. This opacity is approximately that at which CO ice forms in the cloud and it may be that grain growth is initiated via the mediation of the collision damping and sticking characteristics of the ice.
@@
Estimation of extinctions and polarization to understand the polarization efficiency in starless cores. The degree of polarization is found declining beyond extinctions (Av) of 20. This phenomena is named as ’polarization hole’ in molecular cloud cores specially the ones without internal sources a.k.a. starless/prestellar cores such as L183 (Soam & Andersson et al., in prep.) and cloud 109 of the Pipe nebula. Optical spectroscopy of background stars in many such cores are needed to study the variation of polarization with extinction/column density in the outer parts where the column density is too low to allow accurate far infrared observations. That will increase the sample and provide statistical significant results ’polarization hole’ and help understand the critical densities beyond which this happens. Polarization efficiencies i.e. a ratio of degree of polarization and extinction (Av) will also be investigated in a number of starless cores to understand the dust grain alignment.
@@
Dust properties/distribution in high-latitude clouds and high velocity clouds Dust properties and total-to-selective ratio (R) of high-latitude clouds are found to be different from the regions in the Galactic plane. (ref. 28) found that the polarization efficiency in some such clouds can exceed the ”canonical value” of p/A, by a factor of 2-3. Such high polarization efficiencies are difficult to reconcile with standard assumptions about interstellar grains. Dust properties such as size can constrained by performing spectropolarimetry and fitting Serkowski curve to the background stars shining through the diffuse structures seen at high-galactic latitudes. The maximum polarization can be used as an accurate measure of the alignment efficiency.
@@ 2.3 Extra-galactic studies
Spectropolarimetry of type Ia supernovae In a study by (ref. 29), the attempt was made to investigate the properties of extra-galactic dust by spectropolarimetry of four Type Ia Supernovae in other Galaxies. The anomalous extinction curves of these Supernovae were analyzed to understand the environment in which they occurred. The authors found an exceptionally low Rv value and an anomalous behavior of the continuum polarization curve. Type Ia supernovae are very luminous and have a negligible intrinsic continuum polarization. One possible explanation of the low R values and the progression of polarization curves, with successively smaller values is that the RATD dust destruction mechanism is changing the host-galaxy dust distribution responsible for the polarization. Hence, tracking the polarization curve evolution of bright supernovae over the first few weeks after explosion could provide unique information on the dust characteristics and amount in the host galaxies.
@@ 3 Telescope and Instruments
Most sources, especially stars, emit in optical wavelengths with low polarization signals of 1 % or less in . Thus, polarimeters need to be designed to achieve accuracies of 0.1 % or better.
Telescope and optical elements inside an instrument introduce polarization, which leads to the polarization of the beam getting modulated and modified. To estimate the true polarization of a source with accuracies of 0.1 % or better, the polarization behaviour of the instrument has to be carefully modelled and corrected during operational lifetime. Most often, for narrow FoV polarimeters, this is done by observing “polarimetric standard” stars on sky. For wide FoV instruments, dedicated and more involved calibration methods are required.
@@
Most polarimeters built till date have narrow FoV of a few arcminutes or less, mostly allowing for point source polarimetry. New generation polarimeters are being built which will have capability of wide-field imaging polarimetry such as Wide-Area Linear Optical Polarimeters (WALOPs) for the PASIPHAE survey. The design and calibration of our proposed spectropolarimeter will be based on the methods developed for the WALOP instruments. The top level concept of WALOP instrument is shown in Figure 1, taken from (ref. 20). WALOPs employ the concept of four-channel and four-camera one-shot polarimetry, whereby each of the polarized beams (channels) coming out of the polarization analyzer system is steered and imaged into its own camera detector. This technique has the following three major advantages over conventional polarimetry:
@@
Polarimeters in astronomy usually employ two-channel polarimetry, which requires multiple exposures to be taken at various orientations of the analyzer system to estimate the two linear Stokes parameters and . During any such observation sequence, changing sky conditions as well moving/rotating optical components can introduce measurement errors and lower the instrument’s accuracy. To overcome this, one-shot polarimetry has been devised that allows both the linear Stokes parameters to be obtained in one-shot, as demonstrated by the success of RoboPol polarimeter.
@@ 3.1 Case for a 4 m telescope
| sky brightness | 20 mags/arcseconds in V band |
@@ 3.2 Possible instrument design concepts
Polarization gratings: We can combine dispersion and polarization analyzer systems into one by using polarization gratings. As the name may indicate, these novel and upcoming gratings have the combined capabilities of separating the two circular polarizations while also acting as a dispersion element; these elements can provide an accuracy of better than 0.1% . As these separate the two circular polarization components, we will need a rotating quarter-wave plate upstream to measure linear polarization.
@@
Calcite Wollaston prisms: We can use readily available Wollaston prisms made of Calcite for a desired resolution of approximately 50. This approach is being attempted with the RSS instrument on SALT telescope.
@@
Single/Double Wollaston Prisms. These are used ubiquitously in astronomy to obtain high accuracy polarimetric measurements.
@@
Wire-grid polarization beam-splitter: they have uniform behavior over large wavelengths. For example, the MOPTOP instrument uses this for imaging polarimetry. The choice of wire-grid may provide an accuracy of 0.2% on the sky.

### v2

@@ 1 Introduction
Interstellar magnetic fields can be mapped with the polarization of starlight caused by aligned dust grains in different lines-of-sights and in different environments. This was first observed in optical wavelengths11,12 and was explained to be arising from dichroic extinction by elongated dust grains aligned with the ambient magnetic field. Interstellar polarization also provides an opportunity to study the dust grain characteristics e.g. size distribution and composition.
An important science to be addressed with a spectropolarimeter is the investigation of interstellar dust. The differential extinction giving rise to the polarization curve (parameterized by the empirical, convex, “Serkowski function”36
where is the maximum polarization, located at , and “K” controls the width of the curve. As shown by (ref. 17), the Serkowski form can be understood in terms of Mie scattering off an underlying MRN24 grain size distribution - if grains smaller than 0.045m are not aligned. The Serkowski fit to the polarization spectrum of star-light shinning through this interstellar dust can therefore constrain the size of aligned dust grains17. Adding extinction information to the amount of polarization provides important complementary information that can break the degeneracy between the underlying grain size distribution, and the size dependent grain alignment mechanisms, and their efficiencies. Combining spectroscopy and polarimetery is a time efficient and better approach to perform such investigations in comparison to the often used technique of imaging polarimetry with use of filters.
@@ 2.1 Interstellar Medium Science
Spectropolarimetry of stars in or behind nearby star-forming regions: This type of study is needed to obtain the polarization curve for a number of environments and source types. This provides a direct means to probe the size distribution of ISM dust2. It is a powerful technique that for general ISM studies need 2-4 m telescopes, which can use background stars of any spectral class and provides inherently complementary probes to infra-red data. For spatially restricted regions, where the background star selection is severely restricted, requiring a fainter magnitude limit, 4 m telescopes become necessary. As discussed above, the ISM polarization spectrum shows a convex shape with a peak in the range = 0.40.7m35,41. Modern “Radiative Alignment Torque (RAT) alignment” theory17, explains this relation by predicting that grains are aligned if exposed to an anisotropic radiation field with wavelengths less than the grain diameter. A universal linear relationship between and extinction () has been found by (ref. 3) .
(2) where is the average total-to-selective extinction. This result can be understood as due to the progressive reddening of the radiation field into the clouds, but with a fixed, underlying, total grain size distribution throughout each cloud. The dependence on indicates that the average grain size varies from cloud to cloud. As discussed below in item 4, grain growth in the denser parts of the cloud can be traced41 from the deviations of the linear relationship in equation 2.
@@
Investigating grain alignment efficiencies in the regions filled with FUV radiation (high intensity radiation fields) This project is an extension of our study of the grain alignment variations around the wall of the Local Bubble25 and our recently published paper on IC59 and IC63 clouds illuminated by B0 IVe star Cas38. We used our polarization and extinction measurements to estimate polarization efficiencies in these clouds. We investigated how these polarization efficiencies vary with the flux from gamma Cas. Is this UV illumination related to the polarization and hence grain alignment efficiencies in the clouds? This experiment was performed on IC 59 and IC 63 and we compared our results with other regions. We saw a change in the slope of the in polarization efficiencies as we went toward the higher illumination sides. Since IC 59/63 clouds are relatively closer to their illuminating source as opposed to regions like local bubble, we saw high poloarization efficiencies in these clouds. Figure 1 below shows these results. However, as we move towards the regions with more luminous source(s), such as LDN 2046 and especially the Orion Veil, we see that the polarization efficiency seem to saturate (Soam & Andersson., in prep.). We speculate that this is either because in the clouds closer to the radiation source all the alignable grains have aligned by the radiation, or that as the radiation field becomes very strong the large grains are destroyed through the so called RATD destruction15. Because both RAT alignment and RAD Destruction are more efficient for larger grains, but change the polarization efficiency in opposite directions, broad-band spectro-polarimetry would allow the two possibilities to be separated. Our proposed spectropolarimeter is the best suit for that purpose.
@@
The circumstellar envelopes of asymptotic giant branch (AGB) stars provide unique mineralogical laboratories for studying the effects on the grain alignment. This is because of the dredge-up of newly synthesized carbon during the thermal pulses of the AGB phase and the stability of the CO molecule. When the star first ascends the AGB the circumstellar envelope (CSE) is oxygen rich with all the carbon tied up in CO molecules. The dust is therefore made up of silicate and metal oxide grains. As the dredge-up of newly processed material is transferred into the CSE, the C/O ratio eventually exceeds unity and all the oxygen is tied up in CO, with the dust now made up from carbon materials. Thus, the two types of AGB envelopes provide the opportunity to study the silicate and carbon dust separate from each other. As with interstellar medium studies, the Serkowski-curve parameters can yield important information about the dust in the CSE. Grain formation and its influences on the formation and characteristics of the AGB winds can be studied by probing the light from field stars background to the AGB CSE. Models of the initiation and mass loss characteristics of these winds are still an unresolved issue. The brightest AGB CSEs can be probed by FIR and sub-mm wave data at high angular resolution. However, for such emission polarimetry the complex interaction of dust temperature, emissivity and column density/opacity can be difficult to disentangle. With telescopes and polarimeters capable of reaching V 16 in (at least) BVRI polarization curves at several different points in the CSEs can be acquired, probing the alignment efficiency as well as the grain size distribution (at different radii). In initial studies of the IRC+10216 CSE we see a large and K (although the Serkowski form may not be ideal) indicating large grains as compared to the ISM. This is consistent with the results of ab initio AGB wind models33,4, but better S/N is needed. We note that, in the intense radiation field from the central star, the RATD grain destruction mechanism should be important in these CSEs, Measuring spectral types (and thereby the extinction), and polarization of a number of stars shining through the envelopes of the AGB stars can provide observational constrains on the dust grain size in the AGB stars. The fact that our observations indicate large grains without a small grain population indicates that such destruction is not important. This may indicate grains with high tensile strength and hence compact solid structures. Again more and better data are needed to fully explore these possibilities. Based on FIR/sub-mm wave maps of the nearest AGB stars and Gaia distance and photometry many of the nearby AGB stars have several (in some cases 10) background field stars within the projected 70m extent of their CSEs. An extension of such studies would be to probe the ionized regions of planetary nebulae, where the evolved star and dust interacts with the hard UV of the central star.
@@
Grain Growth and Destruction The polarization curve, under RAT alignment, is determined by the size distribution of aligned grains and the SED of the radiation field. The smallest aligned grain is given by the largest of the smallest grain that satisfies the RAT criterion and smallest grain not disaligned by collisions41.
d, where is the shortest wavelength available and d is the grain diameter: Lyman limit (912Å) for the diffuse, neutral, ISM; shorter for fully ionized gas close to hot stars7, or longer for material inside of ISM clouds where reddening is active.
Thermal collisions disalign the smallest grains first. The collision rate goes at and the grains disalignment time goes as 1/d. Hence for denser, warmer gas the limit of aligned grains grows41. The over-all grain size distribution, and the upper cut-off of it is determined by grain growth and destruction.
In dense gas (clumps in clouds) the density can become large enough that significant grain growth can occur41, extending the size distribution and affecting the red part of the polarization curve.
If significant grain destruction (in ISM shocks, or through the RATD mechanism) take place, the largest grains may be destroyed shifting the polarization curve to the blue9.
These various effects can be probed by high-quality determination of the polarization curve in different environments. (ref. 41) studied the polarization curve variations in the Taurus molecular cloud and found evidence for the effects of reddening, gas-dust collisions disalignment and grain growth in clumps dark and cold enough for ISM ices to have formed. This is manifested as a bifurcation in the vs. AV relation (equ. 2), such that, beyond A mag., a second, much steeper relationship was added. This opacity is approximately that at which CO2 ice forms in the cloud and it may be that grain growth is initiated via the mediation of the collision damping and sticking characteristics of the ice.
@@
Estimation of extinctions and polarization to understand the polarization efficiency in starless cores. The degree of polarization is found declining beyond extinctions (Av) of 20. This phenomena is named as ’polarization hole’ in molecular cloud cores specially the ones without internal sources a.k.a. starless/prestellar cores such as L183 (Soam & Andersson et al., in prep.) 1,16 and cloud 109 of the Pipe nebula1. Optical spectroscopy of background stars in many such cores are needed to study the variation of polarization with extinction/column density in the outer parts where the column density is too low to allow accurate far infrared observations. That will increase the sample and provide statistical significant results ’polarization hole’ and help understand the critical densities beyond which this happens. Polarization efficiencies i.e. a ratio of degree of polarization and extinction (Av) will also be investigated in a number of starless cores to understand the dust grain alignment.
@@
Dust properties/distribution in high-latitude clouds and high velocity clouds Dust properties and total-to-selective ratio (RV) of high-latitude clouds are found to be different from the regions in the Galactic plane. (ref. 28) found that the polarization efficiency in some such clouds can exceed the ”canonical value” of p/A, by a factor of 2-3. Such high polarization efficiencies are difficult to reconcile with standard assumptions about interstellar grains13. Dust properties such as size can constrained by performing spectropolarimetry and fitting Serkowski curve to the background stars shining through the diffuse structures seen at high-galactic latitudes. The maximum polarization can be used as an accurate measure of the alignment efficiency.
@@ 2.3 Extra-galactic studies
Spectropolarimetry of type Ia supernovae In a study by (ref. 29), the attempt was made to investigate the properties of extra-galactic dust by spectropolarimetry of four Type Ia Supernovae in other Galaxies. The anomalous extinction curves of these Supernovae were analyzed to understand the environment in which they occurred. The authors found an exceptionally low Rv value and an anomalous behavior of the continuum polarization curve. Type Ia supernovae are very luminous and have a negligible intrinsic continuum polarization. One possible explanation of the low RV values and the progression of polarization curves, with successively smaller values is that the RATD dust destruction mechanism is changing the host-galaxy dust distribution responsible for the polarization. Hence, tracking the polarization curve evolution of bright supernovae over the first few weeks after explosion could provide unique information on the dust characteristics and amount in the host galaxies.
@@ 3 Telescope and Instruments
Most sources, especially stars, emit in optical wavelengths with low polarization signals of 1 % or less12 in . Thus, polarimeters need to be designed to achieve accuracies of 0.1 % or better.
Telescope and optical elements inside an instrument introduce polarization, which leads to the polarization of the beam getting modulated and modified. To estimate the true polarization of a source with accuracies of 0.1 % or better, the polarization behaviour of the instrument has to be carefully modelled and corrected during operational lifetime. Most often, for narrow FoV polarimeters, this is done by observing “polarimetric standard” stars on sky31,32,5. For wide FoV instruments, dedicated and more involved calibration methods are required10,21,22.
@@
Most polarimeters built till date have narrow FoV of a few arcminutes or less, mostly allowing for point source polarimetry. New generation polarimeters are being built which will have capability of wide-field imaging polarimetry such as Wide-Area Linear Optical Polarimeters (WALOPs)20 for the PASIPHAE survey19,39. The design and calibration of our proposed spectropolarimeter will be based on the methods developed for the WALOP instruments. The top level concept of WALOP instrument is shown in Figure 1, taken from (ref. 20). WALOPs employ the concept of four-channel and four-camera one-shot polarimetry, whereby each of the polarized beams (channels) coming out of the polarization analyzer system is steered and imaged into its own camera detector. This technique has the following three major advantages over conventional polarimetry:
@@
Polarimeters in astronomy usually employ two-channel polarimetry, which requires multiple exposures to be taken at various orientations of the analyzer system to estimate the two linear Stokes parameters and . During any such observation sequence, changing sky conditions as well moving/rotating optical components can introduce measurement errors and lower the instrument’s accuracy. To overcome this, one-shot polarimetry has been devised that allows both the linear Stokes parameters to be obtained in one-shot, as demonstrated by the success of RoboPol polarimeter31.
@@ 3.1 Case for a 4 m telescope
sky brightness | 20 mags/arcseconds2 in V band |
@@ 3.2 Possible instrument design concepts
Polarization gratings: We can combine dispersion and polarization analyzer systems into one by using polarization gratings. As the name may indicate, these novel and upcoming gratings have the combined capabilities of separating the two circular polarizations while also acting as a dispersion element27,40; these elements can provide an accuracy of better than 0.1% . As these separate the two circular polarization components, we will need a rotating quarter-wave plate upstream to measure linear polarization.
@@
Calcite Wollaston prisms: We can use readily available Wollaston prisms made of Calcite for a desired resolution of approximately 50. This approach is being attempted with the RSS instrument on SALT telescope30.
@@
Single/Double Wollaston Prisms. These are used ubiquitously in astronomy to obtain high accuracy polarimetric measurements31,32.
@@
Wire-grid polarization beam-splitter: they have uniform behavior over large wavelengths. For example, the MOPTOP instrument uses this for imaging polarimetry37. The choice of wire-grid may provide an accuracy of 0.2% on the sky.